10. Using sample 0.10, 0.11, 0.11, 0.12, 0.13, 0.12,0.11, 0.12, 0.11, 0.11 test the hypothesis MX = 0.10 against the left-hand alternative MX normality, choose ax-5%. 0.10. Assume
Answer #1

Sample size, n =10 <30, small sample. So, we use t-test.

t-test for the population mean(MX):

Null hypothesis (H0): Mean, MX =0.10

Alternative Hypothesis (H1): Mean, MX<0.10

Sample mean, ar{X} =Sigma Xi/n =1.14/10 =0.114​​​​​​

Standard deviation, s =Σ(Xi-X)2/02-1) =0.0084​​​​​​

Test statistic, t =(X - MX)/(s/Vn) =(0.114 -0.10)/(0.0084/V10) =5.27

At degrees of freedom, df =n - 1 =10 - 1 =9 and alpha =5% for left-tailed test, the critical value of t is: tcrit = -1.833

Conclusion:

We failed to reject the null hypothesis at 5% significance level because the test statistic of 5.27 does not fall in the rejection region which is shown in the following image:

S2 七-1.833

Thus, there is no sufficient evidence to claim MX<0.10